# Linear Combination of Saved Checkpoints Makes Consistency and Diffusion Models Better

### v2

@@ 2.1 Diffusion Probabilistic Model
Diffusion models (DM) [36, 37, 8, 28, 40, 15, 4] undertake the reverse operation by initiating with $\bm{x}_{T}$ sampled from pure Gaussian noise and progressively denoising it to reconstruct the image $\bm{x}_{0}$. Importantly, SDE has its corresponding “probability flow” Ordinary Differential Equation (PF ODE) [40, 37], which delineates a deterministic pathway that yields the same distribution $p_{t}(\bm{x})$ for $\forall t$, thereby offering a more efficient sampling mechanism:
@@
where $\nabla\log p_{t}(\bm{x})$ is referred to as the score function of $p_{t}(\bm{x})$ [10]. Various ODE solvers have been introduced to further expedite the sampling process utilizing Eq. 2 or minimizing the truncation error [23, 14, 15, 24].
@@
2.2 Consistency Models
Drawing inspiration from DM theory, consistency models (CM) have been proposed to enable single-step generation [39, 38]. Whereas DM incrementally denoises an image, e.g., via the PF ODE, CM denoted by $f_{\bm{\theta}}$ is designed to map any point $\bm{x}_{t}$ at any given time $t$ along a PF ODE trajectory directly to the trajectory’s initial point $\bm{x}_{t}$ in a single step.
@@ 2.2 Consistency Models
where $f_{\bm{\theta}^{-}}$ refers to the target model and $\bm{\theta}^{-}$ is computed through EMA (see Eq. 6) of the historical weights of $\bm{\theta}$, $\hat{\bm{x}}_{t_{k}}^{\phi}$ is estimated by the pre-trained diffusion model $\phi$ through one-step denoising based on $\bm{x}_{t_{k+1}}$, and $d$ is a metric implemented by either the $\ell_{2}$ distance or LPIPS [49].
@@
where the target model $f_{\bm{\theta}^{-}}$ is set to be the same as the model $f_{\bm{\theta}}$ in the latest improved version of the consistency training [38], i.e. $\bm{\theta}^{-}=\bm{\theta}$.
2.3 Weight Averaging
The integration of a running average of the weights by stochastic gradient descent (SGD) was initially explored within the realm of convex optimization [34, 30]. This concept was later applied to the training of neural networks [42]. Izmailov et al.[12] suggest that averaging multiple weights over the course of training can yield better generalization than SGD. Wortsman et al.[45] demonstrate the potential of averaging weights that have been fine-tuned with various hyperparameter configurations. The Exponential Moving Average (EMA) employs a specific form that uses the exponential rate $\gamma$ as a smoothing factor:
@@ 2.3 Weight Averaging
Practitioners often opt for advanced optimizers such as Adam [16] for different tasks and network architectures, which might reduce the need for employing a running average. However, the use of EMA has been noted to significantly enhance the quality of generation in early DM studies [40, 4, 28, 8, 37]. This empirical strategy has been adopted in most, if not all, subsequent research endeavors. Consequently, CM have also incorporated this technique, discovering EMA models that perform substantially better.
@@
Recently, several works of more advanced weight averaging strategies have been proposed for Large Language Models (LLMs). Most of them focus on merging models fine-tuned for different downstream tasks to create a new model with multiple capabilities [11, 46, 13, 48]. Different from these approaches, our work aims to accelerate the model convergence and achieve better performance in a standalone training process. Moreover, the methodology we employ for determining averaging coefficients is novel, thereby distinguishing our method from these related works. Further details can be found in Sec. 3 and Sec. 4.
2.4 Search-based Methods for Diffusion Models
@@
Search algorithms are widely used across various domains like Neural Architecture Search (NAS) [32, 29], where they are employed to identify specific targets. Recently, many works studies have set discrete optimization dimensions for DMs and utilized search methods to unearth optimal solutions. For example, Liu et al. [21], Li et al. [19] and Yang et al. [47] use search methods to find the best model schedule for DMs. Liu et al. [22] apply search methods to find appropriate strategies for diffusion solvers. However, these works do not involve modification to model weights during search and are only applicable to DMs.
3 Motivation
(a) Metric Landscape of DM.
(b) Metric Landscape of CM.
Figure 3: The metric landscape of the diffusion model. Selected checkpoints $\theta_{n_{0}}$, $\theta_{n_{1}}$, and $\theta_{n_{2}}$ are aligned sequentially along the training trajectory, with $n_{0}<n_{1}<n_{2}$ denoting the progression in the number of training iterations. The origin point $(0,0)$ corresponds to the checkpoint $\theta_{n_{1}}$, while the X and Y axes quantify the differences between $\theta_{n_{1}}-\theta_{n_{0}}$ and $\theta_{n_{2}}-\theta_{n_{0}}$, respectively. A weight located at coordinate $(x,y)$ is formulated as $\theta_{(x,y)}=\theta_{n_{0}}+x(\theta_{n_{1}}-\theta_{n_{0}})+y(\theta_{n_{2}}-\theta_{n_{0}})$. Additional visualizations are provided in App. 0.A.
One notable challenge associated with DMs and CMs is their difficulty in training. We observe that the objectives for DMs and CMs tend to produce substantial variance in gradient estimations, leading to fluctuations in the quality of generation by the resulting models. For DMs (refer to Eq. 3), even if the networks $\bm{s}_{\bm{\theta}}$ are finely tuned to perfectly mirror the score function $\nabla\log p_{t}(\bm{x})$ for any given $\bm{x}_{t}$, the loss as defined does not reduce to zero. As a result, the batched gradient estimation for this objective exhibits high variance. For CMs (refer to Eqs. 4 and 5), at any given time $t$, the network $f_{\bm{\theta}}$ is solely optimized to align with the target model $f_{\bm{\theta}^{-}}$ from the previous time step, $t-1$. This implies that any loss of image information in the early time steps could result in inaccurate gradient estimation in subsequent time steps.
To explore the training challenges of DMs and CMs, we visualize the landscape of generation quality in a 2-dimension (2D) plane. Specifically, we select 3 checkpoints of the weights $\theta_{n_{0}},\theta_{n_{1}},\theta_{n_{2}}$ from the same training trajectory. We then sweep across the 2D space spanned by these three points to assess the FID scores. The results are shown in the first row of Fig. 3. As we can see, there exist model weights with much higher generation quality than any of $\{\theta_{n_{0}},\theta_{n_{1}},\theta_{n_{2}}\}$ in the sub-space spanned by those three checkpoints for both DM and CM. Additionally, results at different training iterations indicate that this pattern of landscape remains stable in different training stages. These insights imply that there exists a substantial opportunity to enhance the model’s performance at any training phase through the linear combination of existing weight checkpoints.
EMA [42] is simple yet widely used technique in the training of DMs and CMs. It operates by exponentially averaging the weights of all previous checkpoints (see Eq. 6), which stabilizes the training process of DMs and CMs to a certain extent. Nonetheless, EMA represents an empirical method of averaging, suggesting potential areas for enhancement. To investigate this hypothesis, we conducted an experiment where the three checkpoints from previous tests were substituted with three EMA weights, followed by a comprehensive re-sweep of all coordinates. The results are demonstrated in the second row of Fig. 3. Notably, it is observed that a linear combination of the three EMA weights could achieve superior performance compared to each individual weight. It is important to highlight that the linear combination of three EMA weights at various training iterations cannot be replicated by simply adjusting EMA rates. Consequently, this suggests that EMA may indeed be a sub-optimal solution, indicating possibilities for further improvement in this area.
Based on the above insights and the observation, we propose LCSC, standing for Linear Combination of Saved Checkpoints, to enhance the training process. Since the basin in Fig. 3 extends beyond the convex hull of the model weights, we further expand the scope of the combination coefficients to encompass any real values and introduce an optimization method to search for the optimal linear combination of the weights.
4 Method
We discuss LCSC in this section. The search problem is defined in Sec. 4.1. In Sec. 4.2, we elaborate on our algorithm. Finally, we list several typical use cases of our method in Sec. 4.3.
4.1 Problem Definition
Let $\{1,2,\ldots,N\}$ denote the set of training iterations. Suppose $\{n_{1},n_{2},\ldots,n_{K}\}$ is a subsequence selected from $\{1,2,\ldots,N\}$. Define $\Theta$ as the set of checkpoints saved at these specific training iterations: $\Theta=\{\bm{\theta}_{n_{1}},\bm{\theta}_{n_{2}},\bm{\theta}_{n_{3}},\cdots,\bm{\theta}_{n_{K}}\}$. Given $\Theta$, we aim to find a group of coefficients $\bm{\alpha}=\{a_{1},a_{2},a_{3},\cdots,a_{K}\}$. These coefficients are intended to linearly combine all the checkpoints in such a way that the resulting model achieves the optimal performance:
$\mathop{\arg\min}\limits_{\begin{subarray}{c}a_{1},a_{2},a_{3},\cdots,a_{K}\in\mathbb{R}\end{subarray}}\mathcal{F}(a_{1}\bm{\theta}_{n_{1}}+a_{2}\bm{\theta}_{n_{2}}+a_{3}\bm{\theta}_{n_{3}},\cdots,a_{K}\bm{\theta}_{n_{K}});\quad\text{s.t.}\quad\sum_{i=1}^{K}a_{i}=1,$ (8)
where $\mathcal{F}(\bm{\theta})$ denotes an evaluation function that measures the generation quality of $\bm{\theta}$ regarding a specific metric, indicating higher quality with a smaller value. The constraint $\sum_{i=1}^{K}a_{i}=1$ reduces the dimension of our search space by 1. However, we find it enables a more effective exploration for the search algorithm.
4.2 Evolutionary Search
When the number of combined checkpoints $K$ is large (e.g., ranges approximately from 200 to 500 in our experiments), the grid search method used in Sec. 3 becomes inefficient. Consequently, we propose to use evolutionary search  [32, 21, 19, 47, 22] to solve the problem.
Our detailed algorithm can be found in Alg. 1. Given a set of checkpoints $\Theta$, we can apply this algorithm at any training iteration $n$ along the training trajectory. We initialize the population using EMA coefficients corresponding to several different rates. Then we conduct an evolutionary search which views a group of coefficients as an individual. In each search iteration, only the top-performing individuals are selected as parents to reproduce the next generation through crossover and mutation. Specifically, during crossover, we randomly mix the coefficients of the two parents or select one of them. For mutation, we introduce random Gaussian noise to each coefficient. This stochastic element ensures that positive adaptations in the coefficients are conserved and advanced to the subsequent generation, whereas detrimental modifications are discarded. Finally, we normalize the new coefficients to adhere to the regularization condition in Eq. 8. We repeat this reproduction process for a predetermined number of times within each search iteration and update the whole population with all newly generated individuals. Upon the completion of the search process, we choose the best individual as the output of our search algorithm.
@@ 4.2 Evolutionary Search
0:   $P$: The whole Population of model schedule.  $CP$: The Candidate Parents set of each loop, from which a parent model schedule is selected.  $NG$: The Next Generation newly mutated from the parent schedule in each loop.  $\bm{\alpha}$: A group of combination coefficients denoted as $\bm{\alpha}=\{a_{0},a_{1},\cdots,a_{K}\}$  $\bm{\alpha}\cdot\Theta$: Equal to $\sum^{K}_{i=1}\alpha_{i}\bm{\theta}_{n_{i}}$
@@
4:  for $t=1,\cdots,\textbf{Iter}$ do
@@
6:     for $i=1,\cdots,\textbf{Epoch}$ do
@@
10:        $\bm{\alpha}_{new}\leftarrow\bm{\alpha}_{new}/sum(\bm{\alpha}_{new})$
11:        $NG\leftarrow NG\cup\{(\bm{\alpha}_{new},\mathcal{F}(\bm{\alpha}_{new}\cdot\Theta))\}$
12:     end for
13:     $P\leftarrow P\cup NG$
14:  end for
@@
15:  $\bm{\alpha}^{*}\leftarrow\mathop{\arg\min}\limits_{\begin{subarray}{c}\bm{\alpha}\in P\end{subarray}}\mathcal{F}(\bm{\alpha}\cdot\Theta)$
16:  return  $\bm{\alpha}^{*}\cdot\Theta$
Given the absence of neural back-propagation in the search process, LCSC can be viewed as a gradient-free optimization method, akin to methods discussed in [41, 27]. This characteristic renders LCSC highly efficient in terms of computation, offering significant savings in GPU memory usage (applicable to both DM and CM) and computational time (specifically for CM, given that the inference process for DM is notably resource-intensive). Furthermore, LCSC possesses the unique advantage of optimizing models directly towards non-derivable metrics where gradient-based methods falter. It is also important to underscore that the applicability of LCSC extends beyond CM and DMs to various other tasks and different models.
4.3 Use Cases
In this section, we delve into the use cases of our method. (1) Our method can be employed during the initial phases of the training process or in scenarios where the training is conducted with a smaller batch size. This application aims to expedite the training process; (2) At the final stage of training, when the model has essentially converged, our method can be utilized to further enhance the model’s performance. This involves refining the output quality of the already converged model and potentially reducing NFE.
Decrease Training Cost. The consumption of training resources is directly influenced by both the number of iterations and the batch size. Through our investigation, we have discerned that the training process for Diffusion Models (DM) and Consistency Models (CM) can be categorized into two distinct phases, as illustrated in Fig. 4. The initial phase is relatively brief, during which DM and CM rapidly acquire the capability to generate visually satisfactory images. In contrast, the second phase is characterized by a slower optimization of models, focusing on the enhancement of sample quality. Implementing LCSC at the onset of this second phase can facilitate an efficient identification of high-quality basins, potentially yielding performance that is on par with, or even surpasses, that of the final converged model. This strategy effectively reduces the total number of training iterations required. Regarding batch size, SOTA DMs and CMs necessitate training with large batch sizes [15, 4, 39], which places a substantial demand on hardware and computational resources. Decreasing the batch size often leads to worse convergence performance. LCSC can be used for low batch size training and achieve equivalent performance to models trained with full batch size.
Enhance Converged Models. LCSC can also be applied to the checkpoints saved during the final stage of training to enhance the performance of converged models. Furthermore, for users who have access to released pre-trained models, we suggest fine-tuning the model for a few iterations and saving the intermediate checkpoints. Then LCSC can be used to enhance the released model.
5 Experiments
In this section we present our experiments to validate the effectiveness of LCSC. In Sec. 5.1, we introduce the experimental settings. Then we demonstrate our results on CM and DM for the two use cases in Sec. 5.2.1. We finally ablate several important hyper-parameters in our workflow in Sec. 5.3.
5.1 Experimental Setup
For DM, we follow DDIM [37] for the evaluation on CIFAR10 [17] and iDDPM [28] for the evaluation on ImageNet-64 [3]. To improve the inference efficiency, we adopt DPM-Solver [24]. For CM [39], we evaluate LCSC with both CD and CT on CIFAR-10 and ImageNet-64. We train models with our own implementation. We follow all the settings reported by the official paper except ImageNet-64, on which we decrease the batch size to 256 on CM and 512 on DM due to the limited resources. We applied LCSC at different training stages. Specifically, for any selected iteration, we utilize checkpoints from every predetermined interval of iterations within a defined window size. Then we run a search process to find the optimal combination coefficients.
Evaluation. We choose the most commonly used metric, FID [7], as the search objective. To evaluate the combination coefficients during the evolutionary search, we sample 5K images on ImageNet-64 with DM and 10K in other settings, using a fixed group of initial noise. For the final model evaluation, we generate 50K samples using another group of initial noises. To demonstrate that LCSC is not overfitted to FID but achieves overall improvement, we report on FID, IS [35], precision [18], and recall [18]. In Sec. 0.B.5, we further report FID@Clip to show that LCSC also brings improvements in other feature spaces.
Baselines. Since EMA is the default setting in almost all works of DM and CM, we report the performance of the EMA weights using the rate reported by official papers [39, 37, 28] as our main baselines. For CM models on ImageNet-64, we additionally download the official models trained with full batch size and test their performance for a complete comparison. We further conduct a grid search of EMA rate for the final model as a stronger baseline, denoted as EMA*, to prove the sub-optimality of EMA and that LCSC is a more effective method. More experimental details are provided in App. 0.B.
5.2 Results
Our results for CM are shown in Tab. 1 and Tab. 2, while for DM, the results are presented in Tab. 3. By conducting LCSC, we consistently find models that outperform the current EMA model across all settings. Although LCSC searches according to FID, consistent improvements are observed across different metrics. Notably, improvement is evident even at the early stages of training, e.g., 50K iterations, when the model is not yet fully trained. Below, we discuss our results under two use cases: decreasing training costs and enhancing the final converged model. Generated images are visualized in App. 0.E.
5.2.1 Results on Decreasing the Training Cost
We conduct a series of experiments using CD and CT to validate that LCSC can significantly reduce training costs. Our reported speedup results have been adjusted to include the search cost on both CPU and GPU, with detailed information available in Sec. 0.B.2.
LCSC can be applied at an early training stage to decrease the number of training iterations. As illustrated in Tab. 1, for CIFAR-10, applying LCSC to the model with only 50K training iterations achieves a better FID than the final model trained with 800K iterations (3.18 vs. 3.65) while accelerating the training by approximately 14× for CD. For CT, we only achieve a converged FID of 9.87 with vanilla training and fail to fully reproduce the result of FID=8.70 reported by Song et al. [39]. However, by applying our method at 400K training iterations, we can still surpass the reported FID by Song et al. [39] with 800K training iterations, achieving around a 1.9× speedup. On ImageNet-64, we decrease the batch size from 2048 to 256 due to limited resources. For CD, LCSC achieves better performance (6.56, 5.71 vs. 7.30 in FID) than the final converged model with only a quarter or half of the training iterations (150K, 300K vs. 600K). For CT, LCSC also significantly reduces the final converged FID (12.05 vs. 15.56) with fewer training iterations (600K vs. 800K).
LCSC can also contribute to reducing the training batch size. On CIFAR-10, we train another model with a lower batch size to validate LCSC’s performance under these conditions. As illustrated in Tab. 1, LCSC with a batch size of 128 outperforms the final converged model with a batch size of 512 for both CD and CT (3.34 vs. 3.65 with CD and 8.33 vs. 8.70, 9.87 with CT), achieving overall speedups of 23× and 7×, respectively. On ImageNet-64, we test the official models of Song et al. [39], which were trained with a 2048 batch size for both CD and CT. With the assistance of LCSC, models trained with a 256 batch size consistently outperform the official models (5.71, 5.54 vs. 6.31 with CD and 12.05 vs. 13.07 with CT), achieving overall acceleration ratios of 15× and 10.4×.
Table 1: Generation quality of consistency models on CIFAR-10. The training speedup is compared against the standard training with 800K iterations and 512 batch size. Our results that beat or match the standard training (the “released” model for FID and IS, and our reproduced results for Prec. and Rec.) are underlined.
CD EMA 50K 512 1 4.84 9.13 0.67 0.55
200K 512 1 4.08 9.18 0.68 0.56
400K 512 1 3.84 9.26 0.68 0.57
800K 512 1 3.66 9.35 0.68 0.57
800K 512 2 2.90 9.56 0.69 0.59
850K 512 1 3.65 9.32 0.68 0.57
850K 512 2 2.89 9.55 0.69 0.58
(released) 800K 512 1 3.55 9.48 - -
(released) 800K 512 2 2.93 9.75 - -
EMA* 800K 512 1 3.51 9.37 0.68 0.57
LCSC 50K 512 1 3.18 9.60 0.67 0.58 $\sim$14$\times$
250K 512 1 2.76 9.71 0.67 0.59 $\sim$3.1$\times$
800K 512 1 2.42 9.76 0.67 0.60 -
800+40K 512 1 2.38 9.70 0.67 0.60 -
100K 128 1 3.34 9.51 0.67 0.57 $\sim$23$\times$
CT EMA 200K 512 1 16.4 8.22 0.64 0.42
400K 512 1 12.1 8.52 0.67 0.43
600K 512 1 10.8 8.67 0.68 0.42
800K 512 1 9.87 8.81 0.69 0.42
(released) 800K 512 1 8.70 8.49 - -
EMA* 800K 512 1 9.70 8.81 0.69 0.42
LCSC 400K 512 1 8.60 8.89 0.67 0.47 $\sim$1.9$\times$
800+40K 128 1 8.05 8.98 0.70 0.45 -
450K 128 1 8.33 8.67 0.69 0.44 $\sim$7$\times$
CD EMA 150K 256 1 8.24 36.2 0.66 0.61
300K 256 1 7.70 37.0 0.67 0.62
600K 256 1 7.30 37.2 0.67 0.62
650K 256 1 7.17 37.7 0.67 0.62
(released) 600K 2048 1 6.31 39.5 0.68 0.63
EMA* 600K 256 1 7.17 37.7 0.67 0.62
LCSC 150K 256 1 6.56 38.9 0.67 0.62 $\sim$29$\times$
300K 256 1 5.71 41.8 0.68 0.62 $\sim$15$\times$
600+20K 256 1 5.54 40.9 0.68 0.62 $\sim$7.6$\times$
CT EMA 200K 256 1 20.3 28.8 0.61 0.53
400K 256 1 17.9 29.8 0.62 0.54
600K 256 1 16.6 30.6 0.62 0.54
800K 256 1 15.8 31.1 0.64 0.55
1000K 256 1 15.6 31.2 0.64 0.55
(released) 800K 2048 1 13.1 29.2 0.70 0.47
EMA* 1000K 256 1 15.6 31.2 0.64 0.55
LCSC 600K 256 1 12.1 35.1 0.67 0.54 $\sim$10.4$\times$
1000K 256 1 11.1 35.7 0.65 0.57 $\sim$6.3$\times$
5.2.2 Results on Enhancing Pre-trained Models
For CM, on CIFAR-10 with CD, CT and ImageNet-64 with CD, to simulate scenarios where users have access only to pre-trained models, we fine-tune the final converged model for a few iterations and then apply LCSC. As illustrated in the 800+40K LCSC row with CD and CT in Tab. 1 and the 600+20K LCSC row with CD in Tab. 2, this approach significantly enhances the performance of the already converged model. For model developers, we directly utilize the checkpoints close to the final iteration, as demonstrated in the 800K LCSC row with CD in Tab. 1 and 1000K LCSC row with CT in Tab. 2. LCSC leads to notable improvements in all these scenarios. Notably, on CIFAR-10, LCSC even outperforms the 800K and 850K models that use 2-step sampling, achieving better performance with 1-step sampling (2.42, 2.38 FID vs. 2.89, 2.90 FID), thereby doubling the inference speed.
For DM, when applying LCSC at the final stage of training, we can identify a model that performs significantly better than the final model obtained through EMA, using the EMA rate selected by the baseline. While a better EMA model may be discovered through an extensive grid search, the outcomes achieved by LCSC consistently surpass those obtained via the EMA strategy, thereby underscoring the superiority of our method. Moreover, we observe that the model identified through our approach can match the performance of the best EMA model with fewer NFE during inference. Specifically, on CIFAR-10, our method requiring NFE=9 is on par with EMA* using NFE=15. Similarly, on ImageNet-64, our method with NFE=12 is competitive with EMA* utilizing NFE=15. This highlights the potential for reducing inference costs while still achieving the baseline’s performance level.
CIFAR10 EMA 50K 15 80.9 5.51 0.52 0.30
150K 15 6.28 8.74 0.61 0.59
350K 15 4.45 9.23 0.63 0.59
800K 15 4.16 9.37 0.64 0.60
EMA* 800K 15 3.96 9.50 0.64 0.60
LCSC 50K 15 10.8 8.35 0.55 0.58
150K 15 4.76 8.97 0.61 0.59
350K 15 3.56 9.35 0.64 0.60
800K 15 3.18 9.59 0.64 0.61
800K 12 3.36 9.56 0.64 0.60
800K 9 3.97 9.50 0.63 0.60
ImageNet EMA 50K 15 59.8 11.0 0.40 0.38
150K 15 22.3 15.0 0.55 0.45
350K 15 20.0 16.4 0.58 0.59
500K 15 19.8 16.9 0.58 0.59
EMA* 500K 15 18.1 17.3 0.59 0.59
LCSC 50K 15 27.2 12.9 0.47 0.56
150K 15 19.1 15.3 0.56 0.56
350K 15 16.8 16.9 0.59 0.59
500K 15 15.3 17.6 0.59 0.59
500K 12 17.2 17.2 0.57 0.59
500K 10 20.5 16.5 0.55 0.58
5.3 Hyperparameter Study
The hyperparameters of LCSC includes the window size ($n_{K}-n_{1}$) for retrieving the historical weights, the interval ($n_{k}-n_{k-1}$) for two adjacent weight checkpoints, the number of samples for computing FID and the number of search iterations. Table 4 llustrates the impact of varying the number of samples, search iterations, and the size of the interval. We observe that increasing the number of samples and search iterations leads to higher performance, though this comes at the expense of increased search cost. In practice, we find that 10K samples and 2K search iterations can efficiently identify strong models. With a fixed window size, the interval between two checkpoints determines the dimension of the search space. Our findings suggest that search performance generally improves as the search space expands, but limiting the search dimension to fewer than 200 can detrimentally affect search performance. A more comprehensive study of the hyperparameters is provided in Sec. 0.B.3.
(a) Results on DM
Samples Search Iters. Interval FID($\downarrow$) IS($\uparrow$)
10K 2K 200 3.87 9.27
2K 2K 200 9.23 9.11
5K 2K 200 4.04 9.09
10K 1K 200 4.01 9.17
10K 4K 200 3.79 9.28
10K 2K 100 3.82 9.15
10K 2K 500 4.41 9.11
(b) Results on CM
Samples Search Iters. Interval FID($\downarrow$) IS($\uparrow$)
10K 2K 100 2.76 9.71
2K 2K 100 3.40 9.39
5K 2K 100 2.79 9.66
10K 1K 100 3.03 9.57
10K 4K 100 2.69 9.70
10K 2K 200 2.84 9.71
10K 2K 50 2.71 9.79
6 Discussion
Figure 5: Visualization of weight combination coefficients obtained using LCSC compared to those from the default EMA on CIFAR10.
In this section, we briefly discuss (1) Search results obtained by LCSC and (2) The reasons why LCSC performs better than traditional SGD.
Analysis of Search Patterns. We visualize several searched combination coefficients in Fig. 5. First, we observe a small subset of weights characterized by coefficients of large magnitude, whereas the coefficients for the majority of weights are nearly zero. This pattern suggests that the basin of the metric landscape likely occupies a subspace of lower rank compared to our entire search space. Additionally, the presence of multiple significant negative coefficients highlights that certain weights act as critical negative examples. This finding further implies that traditional weight averaging methods, such as EMA, are suboptimal, given that their results are confined to the convex hull of the weights, which excludes the discovered basin. More examples illustrating different scenarios are available in Sec. 0.C.1.
LCSC vs.  Graient-based Methods. Our experiments in Sec. 5 have shown the promising benefits of LCSC. Specifically, it can find a basin in the metric landscape based on the weights obtained in early or middle training stages. An intriguing question arises regarding why gradient-based methods do not converge to the basin found by LCSC. In this work, we offer some insights from our research and defer a more comprehensive theoretical investigation to future studies. First, the objective of LCSC is different from gradient-based method. In this work, we employ FID as the search objective as it turns out to be efficient. In principle LCSC can search any objective while gradient-based methods require the objective to be differentiable. It is worth noting that minimizing $\ell_{2}$ loss, which is commonly used in gradient-based optimization, does not necessarily lead to higher generation quality [21, 44]. Second, DM and CM have a large variance in their gradient estimation as discussed in Sec. 3. During training, they may not be able to stay stationary within the basin we have found if the basin is small. A more detailed discussion is provided in Sec. 0.C.1.3.
7 Conclusion
This work proposes a simple but effective method, LCSC, to enhance the training process of DMs and CMs. Based on observations of the metric landscape, we propose to linearly combine saved checkpoints during training to achieve better performance. We further use evolutionary search to find the optimal group of combination coefficients. Our method can be used to decrease the training iteration and batch size, or enhance the final converged models.
Acknowledgement
This work was supported by National Natural Science Foundation of China (No. 62325405, 62104128, U19B2019, U21B2031, 61832007, 62204164), Flemish Government (AI Research Program) and the Research Foundation - Flanders (FWO) through project number G0G2921N, Tsinghua EE Xilinx AI Research Fund, and Beijing National Research Center for Information Science and Technology (BNRist). We thank for all the support from Infinigence-AI.
APPENDIX
Linear Combination of Saved Checkpoints Makes Consistency and Diffusion Models Better
In this appendix, we provide additional examples of the metric landscape in App. 0.A. We further elaborate on our experimental setup and present additional results in App. 0.B. A more in-depth discussion of our work is provided in App. 0.C. Potential directions for future work are explored in App. 0.D. Finally, we offer visualizations of the generated images in App. 0.E. We will open-source our code to promote the application of LCSC.
Appendix 0.A Additional Examples of the Metric Landscape

### v3

@@ B.1 Diffusion Probabilistic Model
Diffusion models (DM) (Sohl-Dickstein et al., 2015; Song et al., 2020a; Ho et al., 2020; Nichol & Dhariwal, 2021; Song et al., 2020b; Karras et al., 2022; Dhariwal & Nichol, 2021) undertake the reverse operation by initiating with $\bm{x}_{T}$ sampled from pure Gaussian noise and progressively denoising it to reconstruct the image $\bm{x}_{0}$. Importantly, SDE has its corresponding “probability flow” Ordinary Differential Equation (PF ODE) (Song et al., 2020b; a), which delineates a deterministic pathway that yields the same distribution $p_{t}(\bm{x})$ for $\forall t$, thereby offering a more efficient sampling mechanism:
@@
where $\nabla\log p_{t}(\bm{x})$ is referred to as the score function of $p_{t}(\bm{x})$ (Hyvärinen, 2005). Various ODE solvers have been introduced to further expedite the sampling process utilizing Eq. 72 or minimizing the truncation error (Liu et al., 2022; Jolicoeur-Martineau et al., 2021; Karras et al., 2022; Lu et al., 2022).
@@
B.2 Consistency Models
Drawing inspiration from DM theory, consistency models (CM) have been proposed to enable single-step generation (Song et al., 2023; Song & Dhariwal, 2024). Whereas DM incrementally denoises an image, e.g., via the PF ODE, CM denoted by $f_{\bm{\theta}}$ is designed to map any point $\bm{x}_{t}$ at any given time $t$ along a PF ODE trajectory directly to the trajectory’s initial point $\bm{x}_{t}$ in a single step.
@@ B.2 Consistency Models
where $f_{\bm{\theta}^{-}}$ refers to the target model and $\bm{\theta}^{-}$ is computed through EMA of the historical weights of $\bm{\theta}$, $\hat{\bm{x}}_{t_{k}}^{\phi}$ is estimated by the pre-trained diffusion model $\phi$ through one-step denoising based on $\bm{x}_{t_{k+1}}$, and $d$ is a metric implemented by either the $\ell_{2}$ distance or LPIPS (Zhang et al., 2018).
@@
where the target model $f_{\bm{\theta}^{-}}$ is set to be the same as the model $f_{\bm{\theta}}$ in the latest improved version of the consistency training (Song & Dhariwal, 2024), i.e. $\bm{\theta}^{-}=\bm{\theta}$.
B.3 Weight Averaging
The integration of a running average of the weights by stochastic gradient descent (SGD) was initially explored within the realm of convex optimization (Ruppert, 1988; Polyak & Juditsky, 1992). This concept was later applied to the training of neural networks (Szegedy et al., 2016). Izmailov et al.(Izmailov et al., 2018) suggest that averaging multiple weights over the course of training can yield better generalization than SGD. Wortsman et al.(Wortsman et al., 2022) demonstrate the potential of averaging weights that have been fine-tuned with various hyperparameter configurations. The Exponential Moving Average (EMA) employs a specific form that uses the exponential rate $\gamma$ as a smoothing factor:
@@ B.3 Weight Averaging
Practitioners often opt for advanced optimizers such as Adam (Kingma & Ba, 2014) for different tasks and network architectures, which might reduce the need for employing a running average. However, the use of EMA has been noted to significantly enhance the quality of generation in early DM studies (Song et al., 2020b; Dhariwal & Nichol, 2021; Nichol & Dhariwal, 2021; Ho et al., 2020; Song et al., 2020a). This empirical strategy has been adopted in most, if not all, subsequent research endeavors. Consequently, CM have also incorporated this technique, discovering EMA models that perform substantially better.
@@
Recently, several works of more advanced weight averaging strategies have been proposed for Large Language Models (LLMs). Most of them focus on merging models fine-tuned for different downstream tasks to create a new model with multiple capabilities (Ilharco et al., 2022; Yadav et al., 2024; Jin et al., 2022; Yu et al., 2023). Different from these approaches, our work aims to accelerate the model convergence and achieve better performance in a standalone training process. Moreover, the methodology we employ for determining averaging coefficients is novel, thereby distinguishing our method from these related works. Further details can be found in Sec. 3 and Sec. 4.
B.4 Search-based Methods for Diffusion Models
@@
Search algorithms are widely used across various domains like Neural Architecture Search (NAS) (Real et al., 2019; Ning et al., 2020), where they are employed to identify specific targets. Recently, many works studies have set discrete optimization dimensions for DMs and utilized search methods to unearth optimal solutions. For example, Liu et al. (Liu et al., 2023a), Li et al. (Li et al., 2023) and Yang et al. (Yang et al., 2023) use search methods to find the best model schedule for DMs. Liu et al. (Liu et al., 2023b) apply search methods to find appropriate strategies for diffusion solvers. However, these works do not involve modification to model weights during search and are only applicable to DMs.
Appendix C LCSC Algorithm
The detailed algorithm of LCSC is provided in Alg. 1.
@@ Appendix C LCSC Algorithm
0:   $P$: The whole Population of model schedule.  $CP$: The Candidate Parents set of each loop, from which a parent coefficients is selected.  $NG$: The Next Generation newly mutated from the parent coefficients in each loop.  $\bm{\alpha}$: A group of combination coefficients denoted as $\bm{\alpha}=\{\alpha_{1},\cdots,\alpha_{K}\}$  $\bm{\alpha}\cdot\Theta$: Equal to $(1-\sum_{i=2}^{K}\alpha_{i})\bm{\theta}_{n_{1}}+\sum^{K}_{i=2}\alpha_{i}\bm{\theta}_{n_{i}}$
@@
4:  for $t=1,\cdots,\textbf{Epoch}$ do
@@
6:     for $i=1,\cdots,\textbf{Iter}$ do
@@
10:        $NG\leftarrow NG\cup\{(\bm{\alpha}_{new},\mathcal{F}(\bm{\alpha}_{new}\cdot\Theta))\}$
11:     end for
12:     $P\leftarrow P\cup NG$
13:  end for
@@
14:  $\bm{\alpha}^{*}\leftarrow\mathop{\arg\min}\limits_{\begin{subarray}{c}\bm{\alpha}\in P\end{subarray}}\mathcal{F}(\bm{\alpha}\cdot\Theta)$
15:  return  $\bm{\alpha}^{*}\cdot\Theta$
Appendix D Additional Examples of the Metric Landscape